OpenStudy (anonymous):

The lengths of the legs of a right triangle are 15 cm and 20 cm. What is the perimeter of the triangle? Note: You will need to use the Pythagorean theorem to find the hypotenuse. A. 35 cm B. 25 cm C. 60 cm D. 300 cm E. 20 cm

Directrix (directrix):

Do you know the Pythagorean theorem?

OpenStudy (phi):

If you memorize 3,4,5 (and multiples of) is a right triangle, this is not too hard

OpenStudy (anonymous):

B. 25 cm

OpenStudy (anonymous):

more than positive

OpenStudy (anonymous):

i can teach you, if you'd like

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

you r big help!!!!!!!!

OpenStudy (anonymous):

...a... big help sorry

OpenStudy (phi):

@nona18 How can the perimeter be less than the sum of the 2 legs?

OpenStudy (anonymous):

sorry, i gave the hypotenuse not perimeter. perimeter is 60 so C.

OpenStudy (phi):

The pythagorean theorem says the longest side of a right triangle (call it c) is found using \[ a^2 + b^2 = c^2 \] You know a and b, so find c then add them up to get the perimeter

Directrix (directrix):

C. 60 cm Continuing from above: 15 + 20 + 25 = 60

OpenStudy (phi):

If you need help on pythagoras, see http://www.khanacademy.org/math/geometry/triangles/v/introduction-to-the-pythagorean-theorem

OpenStudy (anonymous):

were do i go to ask more ?